If the supply is : q = 5p +150 and the demand is q = - 2 + 920 What is the price and the equilibrium quantity ?
1
Expert's answer
2014-03-26T11:21:17-0400
If the supply is : q = 5p +150 and the demand is q = - 2p + 920. What is the price and the equilibrium quantity ?

Solution
The equilibrium price and quantity are in the point where Qd = Qs:
920 - 2P = 150 + 5P;
7P = 770;
P = $110;
Q = 150 +5P = 150+5*110=150+550=700 units.
Answer:$110; 700 units.